OpenStudy (anonymous):

the width of a rectangle is 3 meters less than the length.the perimeter is 14 meters. find the width

OpenStudy (anonymous):

P=2L+2W W=L-3 14=2L+2(L-3) 14=2L+2L-6 14+6=4L-6+6 20=4L 20/4=4L/4 5=L REPLACE P=2L+2W 14=2(5) +2W 14=10+2W 14-10=10-10+2W 4=2W 4/2=2W/2 W=2

OpenStudy (anonymous):

is my answer right